flux quantum (fluxoid) is approximately equal to 2à ƒ ƒ ƒ —10à ƒ ƒ ¢ ˆ ’7 gaussà ƒ ƒ ¢ ˆ ’cm2 A type II superconductor is placed in a small magnetic field, which is then slowly increased till the field starts penetrating the superconductor. The strength of the field at this point is 2à ƒ ƒ  €Ãƒ ƒ ƒ ƒ —105 gauss

The applied field is further increased till superconductivity is completely destroyed. The strength of the field is now 8à ƒ ƒ  €Ãƒ ƒ ƒ ƒ —105gauss. The correlation length of the superconductor is:
a. 20 à ƒ ƒ ƒ …
b. 50 à ƒ ƒ ƒ …
c. 628 à ƒ ƒ ƒ …
d. 2000 à ƒ ƒ ƒ …
1
Expert's answer
2021-10-18T11:12:17-0400

The equation that ties the second critical field and correlation length is


"H_{c2}=\frac{\phi_0}{\pi\epsilon^2}→\epsilon=\sqrt\frac{\phi_0}{\pi H_{c2}}=50\text Å."